OpenStudy (anonymous):

PLEASE HELP!!! A 19.4 gram sample of gas is found to have a volume of 1.3 liters at 282 K and 1.2 atm. What is the molar mass of this gas? Show all of the work used to solve this problem.

OpenStudy (arindameducationusc):

Use PV=nRT

OpenStudy (arindameducationusc):

here T=282 V=1.82 P=1.2 R=0.0821 and n=m/M we need M, so put m=19.4 to get M

OpenStudy (arindameducationusc):

It should look1.2*1.82=(19.4/M)*0.0821*282

OpenStudy (arindameducationusc):

@fenispart Got it?

OpenStudy (anonymous):

@arindameducationusc Yeah I understand now. I got 289.55

OpenStudy (anonymous):

@arindameducationusc I also need help with one last one: A solution is made by dissolving 23.5 grams of glucose (C6H12O6) in .245 kilograms of water. If the molal freezing point constant for water (Kf) is -1.86 °C/m, what is the resulting Δtf of the solution? Show all the steps taken to solve this problem.

OpenStudy (arindameducationusc):

Didn't I give you the formula last time? Okay I will show again...

OpenStudy (arindameducationusc):

\[\Delta{T}=K _{f}* W _{b}*1000/ Mw _{b}*W _{a}\] Here Mwb= 180 because C6H12O6 6*12+12*1+6*1=180

OpenStudy (arindameducationusc):

convert .245 kg to gram....

OpenStudy (anonymous):

25.4g

OpenStudy (arindameducationusc):

you get 245 gram (This is Wa)

OpenStudy (arindameducationusc):

1kg=1000grams

OpenStudy (arindameducationusc):

and Wb is 19.4, and the rest is given in question... Got it?

OpenStudy (arindameducationusc):

@fenispart Can you solve now?

OpenStudy (anonymous):

Is 19.4 always Wb?

OpenStudy (arindameducationusc):

No not always, but for this question 19.4 is Wb

OpenStudy (anonymous):

How do you find Wb?

OpenStudy (arindameducationusc):

hmmmm....Its given in the question.... READ THE QUESTION

OpenStudy (anonymous):

Wait are we talking about the first question or the second one?

OpenStudy (arindameducationusc):

oops... my bad...... its second question and Wb is 23.5.

OpenStudy (arindameducationusc):

sorry about that

OpenStudy (arindameducationusc):

23.5 is given... I guess now you got it. Both questions are about gas so.... I made a mistake........

OpenStudy (arindameducationusc):

Can you do now? @fenispart

OpenStudy (anonymous):

How do I find Kf? @arindameducationusc

OpenStudy (arindameducationusc):

Kf is given -1.86

OpenStudy (anonymous):

I got -0.991 @arindameducationusc

OpenStudy (arindameducationusc):

Total answer?

OpenStudy (anonymous):

Yes

OpenStudy (anonymous):

@arindameducationusc

OpenStudy (arindameducationusc):

i m here

OpenStudy (arindameducationusc):

Yes your answer is correct....

OpenStudy (arindameducationusc):

I just checked it... Good job! Bravo!

OpenStudy (anonymous):

Thank you so much!